OpenStudy (anonymous):

Point P is located at (-2,3) and Point Q is located at (2,7) find the length of segment PQ!

OpenStudy (anonymous):

it's sq rt (32)

OpenStudy (anonymous):

Distance formula is \[D=\sqrt{(x2 - x1)^{2}+(y2-y1)^{2}}\]

OpenStudy (anonymous):

sq rt ( (x2-x1)^2 + (y2-y1) ) = sq rt ( 4^2 + 4^2) = sq rt ( 16+16) = sq rt (32)

OpenStudy (anonymous):

or you could write the final soulution as 4 sq rt(2)